### v2

# Hiding Latencies in Network-Based Image Loading for Deep Learning

Abstract
In the last decades, the computational power of GPUs has grown exponentially, allowing current deep learning (DL) applications to handle increasingly large amounts of data at a progressively higher throughput. However, network and storage latencies cannot decrease at a similar pace due to physical constraints, leading to data stalls, and creating a bottleneck for DL tasks. Additionally, managing vast quantities of data and their associated metadata has proven challenging, hampering and slowing the productivity of data scientists. Moreover, existing data loaders have limited network support, necessitating, for maximum performance, that data be stored on local filesystems close to the GPUs, overloading the storage of computing nodes.
In this paper we propose a strategy, aimed at DL image applications, to address these challenges by: storing data and metadata in fast, scalable NoSQL databases; connecting the databases to state-of-the-art loaders for DL frameworks; enabling high-throughput data loading over high-latency networks through our out-of-order, incremental prefetching techniques. To evaluate our approach, we showcase our implementation and assess its data loading capabilities through local, medium and high-latency (intercontinental) experiments.
Keywords: Data loading, Deep learning, High-Throughput, Latency optimizations, Scalable storage, Image classification.
1 Introduction
Over the last two decades, the rapid increase in GPU computational power has transformed the field of machine learning. This surge in processing capabilities has allowed deep learning (DL) models to manage vast datasets and conduct intricate computations with unmatched efficiency. Consequently, DL techniques have gained widespread traction across various sectors, leading to advancements in areas such as cybersecurity, natural language processing, bioinformatics, robotics and control, and medical information processing, among others [1].
However, despite these advancements in computational power, the performance of DL systems is increasingly constrained by bottlenecks related to data movement and access. While GPUs continue to achieve remarkable gains in processing throughput, which is matched by an increase in bandwidth for networking and storage, the latencies in these areas cannot decrease at the same pace due to inherent physical limitations. These discrepancies result in significant data stalls, as the transfer of data between storage systems, memory, and processing units becomes a critical bottleneck, leading to inefficient resource utilization in DL workflows [14, 19].
Furthermore, managing the vast amounts of data and associated metadata required for DL has become an increasingly complex challenge, significantly impacting the productivity of data scientists. As datasets scale from sources such as ImageNet [5], which comprises millions of images and spans hundreds of gigabytes, to more extensive datasets like LAION-5B [22], containing billions of images and reaching hundreds of terabytes, the exponential growth in data volume places substantial strain on traditional filesystem-based approaches. These systems are often inflexible and ill-suited for dynamic data management requirements. The limitations of these approaches are particularly evident when tasks such as balancing class distributions [6], or adjusting the proportions of training, validation, and test sets, must be performed, especially when considering metadata to avoid introducing unwanted biases. The static nature of traditional filesystems hampers the ability to efficiently update or modify datasets, restricting the flexibility needed for iterative development and fine-tuning of DL models.
Moreover, DL applications often involve datasets comprising numerous small inputs that are fully scanned and randomly permuted at each training epoch. This access pattern diminishes the benefits of caching since the data is continuously reloaded and reshuffled [16]. Additionally, high-performance computing (HPC) storage systems are typically optimized for sequential access to large files, which contrasts sharply with the small, random access patterns typical of DL workloads. As a result, parallel file systems such as GPFS or Lustre in HPC environments struggle to handle these workloads efficiently [21]. A common mitigation strategy is to maintain local copies of datasets on all compute nodes. However, this approach introduces substantial storage overhead and capacity constraints. An alternative is to partition the dataset into shards distributed across nodes, which becomes essential when the dataset exceeds the storage capacity of individual nodes. Yet, ensuring unbiased data sharding poses significant challenges [21]. This strategy also compromises the ability to perform uniform random shuffling, which can negatively impact training performance and model accuracy [25].
To address these limitations, specialized data loading systems and strategies have been developed, as elaborated in the next section. However, these methods have critical shortcomings, leaving unresolved issues such as:
  • Decoupling of data and metadata, leading to potential inconsistencies;
  • Inflexibility of record file formats, which impose constraints on shuffling;
  • Limited support for network-based data loading, resulting in local storage overload on computing nodes.
In response to these issues [24] proposed leveraging scalable NoSQL databases to store both data and metadata, with preliminary performance evaluations conducted within the DeepHealth Toolkit [4], a DL framework tailored for biomedical applications. In this work, we build upon this concept by introducing and evaluating a novel data loader. Specifically, our key contributions are as follows:
  • We develop an efficient data loader implemented in C++ with a Python API, designed to integrate seamlessly with Cassandra-compatible NoSQL databases and NVIDIA DALI [7, 17]. This loader supports data loading across the network and is compatible with popular DL frameworks such as PyTorch and TensorFlow.
  • We introduce out-of-order, incremental prefetching techniques that enable high-throughput data loading, even in high-latency network environments;
  • We conduct a comprehensive evaluation of our approach, demonstrating its implementation and benchmarking its performance through extensive experiments in local, medium and high-latency settings, comparing it against the state-of-the-art tools.
Note that, as DALI is primarily optimized for image processing, our examples will focus on DL applications involving images; however, the techniques presented are of general applicability.
2 Background and related work
In this section, we begin with an overview of record file formats in DL. Next, we review state-of-the-art data loading software, highlighting their advantages and limitations. Finally, we introduce the NoSQL databases that will be utilized by our data loader.
2.1 Record file formats
Many DL applications, such as image classification, exhibit limited temporal and spatial locality due to their scan-and-reshuffle data access patterns [16]. This lack of locality impedes optimization strategies such as block reading, which involves retrieving multiple images in a single request. When adopted, block reading can help mask latency, alleviate stress on the filesystem, and ultimately improve throughput in both storage and network systems. To leverage this optimization, despite the lack of inherent spatial locality, engineers have developed file formats that artificially enforce spatial locality by grouping (unrelated) files together in record files. Examples include TFRecord [20] (initially developed for TensorFlow and also supported by PyTorch [^1] https://pytorch.org/data/main/generated/torchdata.datapipes.iter.TFRecordLoader.html), RecordIO (designed for the now archived MXNet framework [^2] https://attic.apache.org/projects/mxnet.html), Beton (developed within the FFCV project [15]) and MDS (developed within Databricks’s MosaicML platform, see §2.2.4), the first two being also supported by NVIDIA DALI.
However, optimization algorithms like Stochastic Gradient Descent and Adam require uniform shuffling of data for optimal convergence [25]. Block reading conflicts with this requirement, leading to the implementation of workarounds in data loaders. For instance, they may load a window of data several times larger than the desired minibatch size into memory and shuffle internally, ensuring that minibatches vary across epochs. Even so, the resulting shuffle is not uniform, as images stored close together in a record file will always appear in nearby minibatches.
Nevertheless, the primary drawback of the file-batching approach is that it further rigidifies the dataset. Writing record files is time-consuming, consumes additional storage space, and makes it even more challenging to modify datasets – an already cumbersome task when dealing with numerous files in a filesystem.
2.2 State-of-the-art data loaders
2.2.1 NVIDIA DALI
A major inefficiency in a typical naive image classification workflow [^3] E.g., https://pytorch.org/tutorials/beginner/transfer_learning_tutorial.html stems from the fact that image loading and decoding are managed by high-level routines written in Python. Due to Python’s Global Interpreter Lock (GIL), which limits parallel multi-threading, data loading suffers from considerable serialization and copying overhead between processes [18, 12]. Additionally, pre-processing operations such as resizing, rotating, cropping, and normalization are typically performed on the CPU, which can become a bottleneck, causing delays as the faster GPUs remain idle while waiting for data.
NVIDIA DALI is a state-of-the-art data loader [7, 17], which addresses these inefficiencies by managing the entire pipeline of loading, decoding, and pre-processing images. It leverages GPU acceleration for decoding and pre-processing tasks, significantly reducing CPU bottlenecks. DALI integrates seamlessly with both PyTorch and TensorFlow, offering a versatile solution for DL workflows. Its asynchronous, pipelined execution model with prefetching ensures efficient data handling and minimizes idle GPU time. As free software written in C++, with Python bindings, DALI is modular and extensible, allowing users to customize it with their own plugins for specialized tasks.
One current limitation of DALI is its limited support for data loading over networks. As of this writing, it has only recently introduced experimental support for data loading from S3, and its performance remains suboptimal [^4] https://github.com/NVIDIA/DALI/issues/5551. Additionally, DALI does not currently support reading data from more structured sources, such as databases.
2.2.2 TensorFlow’s tf.data
The tf.data module [20] is an efficient framework for data loading and processing, developed as part of the TensorFlow ecosystem. Notably, it is exclusive to TensorFlow and is not supported by PyTorch. Like DALI, it is implemented in C++ for performance reasons but provides a Python API, allowing seamless integration with TensorFlow workflows. One of the key strengths of tf.data is its ability to efficiently manage parallel data loading and preprocessing, significantly reducing input pipeline bottlenecks during training on large datasets. The module supports various data sources, including local file systems and networked storage, facilitating distributed data handling and enabling the construction of scalable and flexible data input pipelines. It offers a range of optimizations, such as prefetching, caching, sharding, and use of record files (TFRecord) to improve overall throughput in DL workflows.
However, its native support for network-based data loading is limited to Google Cloud services, which may pose constraints for users relying on alternative cloud platforms or custom data storage solutions.
2.2.3 tf.data service
The tf.data service [2], still experimental, offers an alternative approach for data loading in TensorFlow. This service extends the tf.data API by decoupling data loading and preprocessing from model training, thus enabling efficient scaling of data input pipelines across multiple workers. By offloading data preprocessing and distribution to a centralized service, hosted on dedicated nodes, the tf.data service aims at enhancing parallel data loading and transformation, helping to alleviate input pipeline bottlenecks. It is worth noting that this service can also be utilized to enable network-based data loading, with the dispatcher and workers supplying data to the nodes responsible for model training.
2.2.4 MosaicML Streaming Dataset
MosaicML Streaming Dataset (SD) library [23] is a Python package designed for efficient and scalable handling of large datasets, especially in machine learning and data processing workflows. It supports datasets that exceed the capacity of a single node by using streaming techniques to load and locally buffer data dynamically during training or evaluation. Key features include the ability to mix multiple data sources with configurable weighting, flexible handling of diverse data types (e.g., images, text, video, and multimodal data), and true determinism to facilitate debugging and reproducibility.
Note that to utilize SD, data must be converted into a proprietary record format (known as Mosaic Data Shard). This format organizes data samples into large files called shards, accompanied by an index file containing metadata. The metadata enable efficient partitioning of samples across nodes and GPUs prior to training, reducing redundant downloads and enhancing performance.
2.2.5 Deep Lake
Deep Lake [9] is a specialized database optimized for DL applications, combining the features of data lakes and vector databases to efficiently manage a wide range of data types, including text, audio, video, images, PDFs, embeddings, and annotations. It is particularly suited for the development of Large Language Model applications and the training of DL models. Deep Lake utilizes a proprietary record file format, Tensor Storage Format, and supports remote data access through AWS S3, Google Cloud Storage, Azure, or its own Activeloop Storage. The platform offers serverless, scalable storage solutions, enabling users to manage diverse datasets in a unified environment.
A key limitation of Deep Lake is that its high-performance C++ data loader for distributed training at scale is closed-source and only available in the premium enterprise version, which requires a paid subscription. Additionally, this faster data loader does not support self-hosting, forcing users to rely on external cloud storage services. This limitation may pose challenges for organizations with specific data residency requirements or those seeking to minimize data transfers outside their internal networks.
2.2.6 MADlib
MADlib [10] is an open-source, in-database machine learning library originally developed for scalable and parallelized machine learning on relational databases. Its core concept is to bring machine learning closer to where the data resides, enabling analytics and training operations to be performed directly inside the database. This approach minimizes the need for large-scale data transfers and leverages the parallel processing capabilities of modern relational database management systems (RDBMS). Initially designed for traditional machine learning algorithms like logistic regression, decision trees, and k-means clustering, MADlib has evolved to support more advanced analytics. Recent developments have integrated support for DL frameworks like Keras, allowing the use of neural networks alongside conventional ML algorithms. However, DL support in MADlib is still at an early stage and has several limitations:
  • Images are stored as uncompressed tensors;
  • Only image classification is currently supported (no segmentation);
  • Minibatches are solidified, i.e., they are saved together as a row in the DB.
Interestingly, MADlib’s trajectory can be interpreted as an inverse approach compared to the methodology examined in our work. While MADlib started as an RDBMS-centric solution for machine learning and is gradually expanding its capabilities toward DL, this work connects traditional DL data loaders to NoSQL database for image storage and retrieval over the network. This highlights a convergence where both approaches – extending database-centric tools to support DL, and using specialized databases to manage the growing complexity of DL data – meet similar challenges from opposite directions. The shared goal is optimizing data management and processing efficiency in large-scale, data-intensive learning scenarios.
2.3 Cassandra-compatible databases
Our data loader, outlined in the following section, retrieves data from Cassandra-compatible databases, which are summarized below.
2.3.1 Apache Cassandra
Apache Cassandra is a secure, distributed, and decentralized NoSQL database system known for its high scalability and fault tolerance. It is specifically designed to support geographically distributed deployments across multiple data centers, ensuring high availability and resilience. Written in Java and open-source, Cassandra is optimized for low-latency operations, typically delivering response times in the range of single to low double-digit milliseconds for small data transactions. These characteristics have made it a popular choice in the big data industry, with notable adopters including Netflix and Spotify.
2.3.2 ScyllaDB
ScyllaDB is a real-time, big-data NoSQL database engineered to be API-compatible with Cassandra. Developed as open-source software in C++, it offers significant performance enhancements over Cassandra, particularly in terms of reduced latency and lower variability. This improvement is largely due to the absence of a Java garbage collector and the use of a shard-per-core architecture. These optimizations make ScyllaDB well-suited for high-performance workloads, as also proven by its adoption by major companies such as Discord, Ticketmaster, and Rakuten.
3 High-performance network data loading: design principles and implementation
The design of our data loader was guided by the following critical objectives, each aimed at addressing the unique challenges posed by modern deep learning workflows:
  • Integration of data and metadata storage: By tightly coupling data with its associated metadata, we minimize the risk of inconsistencies arising from discrepancies between the two. This design choice ensures that datasets remain reliable and self-contained, eliminating potential errors during dataset update.
  • Provision of fast, flexible, and scalable network access: Our architecture supports high-throughput access to data, ensuring that input pipelines do not become a bottleneck in training. It is designed to scale seamlessly across a range of network configurations, from local clusters to cloud-based environments, while maintaining low latency and robust performance.
  • Ensuring unrestricted random access: The ability to access any data sample at random is crucial for shuffling datasets uniformly during training. Our loader is built to guarantee this capability, even when handling large-scale datasets stored across distributed systems, ensuring that training remains unbiased and effective.
  • Simplifying data partitioning and data distribution in parallel settings: By separating storage management from dataset partitioning, our design offers flexibility for dynamic splitting strategies like cross-validation while maintaining a consistent retrieval interface. It also streamlines data distribution across nodes or GPUs, reducing overhead and ensuring efficient operation in distributed training setups.
Together, these design principles underpin a robust, efficient, and user-friendly data loading solution tailored for the demands of DL applications, particularly those involving large-scale image classification tasks. To implement the design principles outlined above, we developed our data loader as a plugin that bridges NVIDIA DALI and a Cassandra-compatible NoSQL database. This approach enables high-performance data retrieval, ensuring scalability and reliability for modern DL workloads. By implementing the data loader in C++, we eliminate the interprocess communication overhead typically associated with Python-based loaders, providing significant improvements in data throughput and latency.
3.1 Data flow and model
The data flow of DL applications utilizing our data loader is depicted in Fig. 1 and proceeds as follows:
  • Data are extracted from the original dataset and stored, along with associated metadata, in a Cassandra-compatible database;
  • Each image is uniquely identified using a Universally Unique Identifier (UUID);
  • Splits, represented as lists of UUIDs, can be automatically generated based on target values and constraints defined by the metadata;
  • When needed, data are efficiently retrieved using their UUIDs, passed to the DALI pipeline for preprocessing, and then fed into the DL engine (either PyTorch or TensorFlow).
Figure 1: High-level architecture of our data loader.
This architecture enables data loading over the network using TCP, allowing full random access to datasets. Additionally, by leveraging Cassandra or ScyllaDB for storage, it offers significant advantages such as easy scalability and secure data access through SSL. It also allows for the definition of roles with detailed permissions and supports straightforward geographic replication.
To optimize performance and simplify queries, we store data and metadata in separate tables. This separation improves retrieval efficiency: metadata, which is solely used for creating data splits, can be queried independently of the data accessed during the training process. However, in order to ensure database consistency, data and metadata associated with the same image are inserted atomically using Cassandra’s BatchStatement.
As an example of image classification that involves a complex set of metadata, we present SQL tables in Listing 1 that are applicable to medical imaging, specifically for tumor detection in digital pathology [3]. In this context, gigapixel images are divided into small patches, with each patch being labeled to indicate the degree or severity of cancerous tissue (Gleason score).
It is important to note that our data loader is designed with flexibility beyond standard image classification tasks. It accepts features stored as a generic BLOB, allowing for any file format decodable by DALI (e.g., JPEG, JPEG 2000, TIFF, PNG, etc.), which also offers the option to include custom decoders if needed. Annotations are optional and can be provided as integer values (for classification) or as another BLOB, enabling support for tasks like multilabel classification (where labels can be stored as serialized NumPy tensors) and semantic segmentation (where masks are stored as images, for instance, in PNG format).
Listing 1 Example of SQL data model for tumor detection
CREATE TABLE patches.metadata(
patient_id text,
slide_num int, // patients can have
// several slides
x int, // coordinates
y int, // within the slide
label int, // Gleason score
patch_id uuid,
PRIMARY KEY ((patch_id))
CREATE TABLE patches.data(
patch_id uuid,
label int, // Gleason score
data blob, // image/tensor file
// (JPEG, TIFF, NPY, etc.)
PRIMARY KEY ((patch_id))
3.2 Automatic Split Creation
Creating dataset splits, such as training, validation, and test sets, requires careful consideration of metadata to ensure both data independence and balance. A key consideration in this process is maintaining entity independence, where specific entities – such as individuals, groups, or sessions – must be assigned to distinct splits to prevent data leakage and preserve the integrity of model evaluation. For instance, in medical datasets, it is crucial to assign patients to separate splits to avoid bias and ensure generalizability. Additionally, it is common practice to adjust class distributions within the splits to reflect desired proportions of each class, which may involve modifying class weights to address imbalances. However, managing these tasks with standard workflows, i.e., including the reorganization of subdirectories or the recreation of TFRecord files, can be highly labor-intensive and prone to error, underscoring the need for automated solutions to streamline the process.
Our plugin automates both the creation of dataset splits and the corresponding data loading process, decoupling storage from splits and eliminating the need for manual creation of TFRecord files or subdirectory reorganization. By automating both split creation and data loading, it enhances workflow efficiency, enabling users to focus on model development without the complexities of managing dataset splits or file organization. This automation not only simplifies the integration process but also improves reproducibility, facilitating consistent and efficient handling of complex datasets.
3.3 Multi-threaded asynchronous data loading
To optimize data retrieval efficiency, we leverage extensive parallelization. Images are retrieved asynchronously across multiple threads and TCP connections, thereby minimizing overall latency. Each TCP connection can handle up to 1024 concurrent requests, with the number of connections being a tunable parameter. Once the images are retrieved, batches of data are assembled in shared memory, which eliminates the need for additional copying and accelerates the process.
In detail, our batch loading workflow starts with the batch loader receiving a list of UUIDs, which it then uses to send all requests to the Cassandra driver at once. Communications for different batches are handled concurrently via a thread pool. To manage these requests efficiently, multiple low-level I/O threads are employed, each utilizing two TCP connections. Results are processed through callbacks, which minimizes latency by eliminating busy waiting. After all results for a batch have been received, the output tensor is allocated contiguously in a single operation. Data is then copied into the output tensor concurrently, again using a thread pool. The batch becomes available for output as soon as the copying is complete. The call graph illustrating this workflow is shown in Fig. 2.
Figure 2: Call graph of our data loader.
3.4 Prefetching techniques and strategies for high-latency environments
Efficient data loading is essential for optimizing training throughput, particularly in high-latency network environments. A common approach leverages the fact that, even if the dataset is shuffled at the start of each epoch, the permutation is predetermined and all the future requests are known at the beginning of each epoch. This allows to apply prefetching techniques, enabling subsequent batches to be retrieved while the GPUs process the current one, thereby minimizing idle time and improving overall efficiency.
Our data loader features a prefetching mechanism with a configurable number of batch buffers, designed to mask latencies of varying magnitudes. Despite this, during preliminary tests over real high-latency internet connections, we observed significant underperformance compared to internal tests with artificially induced latency using tc-netem [11]. Traffic analysis indicated significant bandwidth variability due to multiple TCP connections traversing different network routes, some of which experienced congestion, resulting in a wide disparity between the best and worst-performing connections. This directly impacts batch loading times: since images are retrieved in parallel but assembled in order, the system must wait for the slowest connection before proceeding, which ultimately slows down the entire process.
To address this issue, we implemented an out-of-order prefetching strategy. Given that DL training is robust to uniformly random permutations of the dataset, we can concurrently request multiple batches (e.g., 8) and reassemble them based on the arrival time of the contained images. This approach reduces the impact of slow connections by prioritizing images that arrive first. For this strategy to function correctly, it is essential that labels are retrieved together with their corresponding features, a requirement met by our architecture, which retrieves both features and annotations with a single query.
Further testing over real high-latency internet connections revealed another, second-order, issue: an aggressive filling of the prefetch buffers (e.g., 8 buffers per GPU across 8 GPUs) can cause a burst of requests that temporaly overwhelms the network capacity, leading to unstable throughput during buffer filling. To mitigate this, we can stagger the prefetch requests over time. For example, instead of front-loading all prefetch requests, we can request an extra batch every four regular ones: for every four batches consumed, five new ones are requested until the buffer is full. This approach limits the increase in transient throughput to only 25% above the steady-state level.
These optimizations collectively enhance throughput and stabilize data loading in high-latency environments, significantly improving resource utilization, as detailed in Sec. 4.
4 Software evaluation and results
4.1 Availability and usage
Our data loader is free software released under the Apache-2.0 license and is accessible on GitHub at the following URL: https://github.com/crs4/cassandra-dali-plugin/. The repository includes a Docker container that provides a pre-configured environment with DALI, Cassandra DB, a sample dataset for experimentation, and comprehensive instructions for conducting further tests. The repository provides scripts to streamline the ingestion of datasets into Cassandra, either serially or in parallel via Apache Spark. It also includes examples of multi-GPU DL training in PyTorch, using both plain PyTorch and PyTorch Lightning. Additionally, it features tools for automatic dataset splitting and high-performance inference using NVIDIA Triton [^5] https://developer.nvidia.com/triton-inference-server. This setup enables clients to request inference of images stored on a remote Cassandra server to be processed on a different GPU-powered remote server.
Listing 2 presents the Python code used to initialize a typical DALI pipeline, which includes data loading via the standard file reader, decoding, and standard preprocessing steps such as resizing, cropping, and normalization. To use our custom data loader, which supports data reading from a Cassandra-compatible DB, one simply needs to replace the standard file reader with our module, as demonstrated in Listing 3. The modified lines are highlighted in blue.
Listing 2 Initializing DALI pipeline using DALI standard file reader
@pipeline_def(batch_size=128, num_threads=4, device_id=device_id)
def get_dali_pipeline():
images, labels = fn.readers.file(name="Reader",
file_root="/data/imagenet/train")
labels = labels.gpu()
images = fn.decoders.image(images, device="mixed",
output_type=types.RGB)
images = fn.resize(images, resize_x=256, resize_y=256)
images = fn.crop_mirror_normalize(images, dtype=types.FLOAT,
output_layout="CHW", crop=(224, 224),
mean=[0.485 * 255, 0.456 * 255, 0.406 * 255],
std=[0.229 * 255, 0.224 * 255, 0.225 * 255],
return images, labels
Listing 3 Initializing DALI using our Cassandra-DALI plugin
uuids = list_manager.get_list_of_uuids(…)
@pipeline_def(batch_size=128, num_threads=4, device_id=device_id)
def get_dali_pipeline():
images, labels = fn.crs4.cassandra(name="Reader",
cassandra_ips=[1.2.3.4, 5.6.7.8],
username="guest", password="test",
table=imagenet.data_train, uuids=uuids,
prefetch_buffers=16, io_threads=8
labels = labels.gpu()
images = fn.decoders.image(images, device="mixed",
output_type=types.RGB)
images = fn.resize(images, resize_x=256, resize_y=256)
images = fn.crop_mirror_normalize(images, dtype=types.FLOAT,
output_layout="CHW", crop=(224, 224),
mean=[0.485 * 255, 0.456 * 255, 0.406 * 255],
std=[0.229 * 255, 0.224 * 255, 0.225 * 255],
return images, labels
4.2 Comparative analysis at varying latencies
We evaluated our network data loader alongside two state-of-the-art competitors, leveraging Amazon EC2 instances located in different geographical regions to introduce varying latencies. Specifically, we tested data consumption in Oregon using an 8-GPU p4d.24xlarge node, while storing images at locations characterized by the following latencies:
  • Low: Data stored in Oregon, round-trip time (RTT) < 1 ms.
  • Medium: Data stored in Northern California, RTT $\simeq$ 20 ms.
  • High: Data stored in Stockholm, Sweden, RTT $\simeq$ 150 ms.
While storing images on a different continent from the GPUs is not a common practice, the high-latency scenario was included to highlight the challenges posed by latency-induced bottlenecks. Such challenges are expected to further intensify in the future, as computational power and bandwidth improve, whereas latency remains constrained by physical limits.
For our experiments, we utilized the standard ImageNet-1k dataset, described in Table 1. The dataset was prepared in the formats required by each data loader and stored on a single node equipped with four NVMe SSDs, configured as a single striped logical volume for optimized data access. Specifically, we used r5dn.24xlarge instances in Oregon and Stockholm, and the similar m6in.24xlarge instance in Northern California, where the previous instance type was not available. As the RAM capacity of these machines surpasses the size of our test dataset (i.e., ImageNet-1k), we reserved memory to maintain approximately only 70 GB of free RAM (i.e., half the size of the dataset). This approach prevents dataset caching in main memory, ensuring that data is consistently read from the disks during testing. By doing so, we simulate conditions involving larger datasets that exceed the available memory capacity.
Table 1: ImageNet-1k dataset properties and test parameters
Training set 1,281,167 images
Average image size 115 kB
Total size 147 GB
Batch size 512 images
Table 2: Overview of the tested data loaders.
Data Loader Version Data Storage Granularity
Cassandra-DALI 1.2.0 ScyllaDB (Cassandra-compatible) Single image
MosaicML SD 0.10.0 MinIO S3-compatible server Record file (MDS)
tf.data service 2.16.1 Filesystem on remote node Record file (TFRecord)
Three data loaders, summarized in Table 2, were compared in this study: our Cassandra-DALI data loader (with data stored in high-performance ScyllaDB), MosaicML SD (with data hosted on an S3 MinIO server), and TensorFlow’s tf.data service (with data stored as TFRecords in the filesystem). All servers were hosted on the same node within Docker containers, with all data residing on the same logical volume. The test code, including the Dockerfiles, is available in the following GitHub repository, under the paper branch: https://github.com/fversaci/cassandra-dali-plugin/.
We conducted two experiments to evaluate performance under varying latencies:
  • Tight-loop read: This benchmark assesses raw data-loading capabilities by maximizing data reads without GPU processing or image decoding.
  • Training: A standard PyTorch multi-GPU ResNet-50 training workload, including image decoding and preprocessing steps such as resizing, normalization, and cropping.
It is important to note that the tight-loop read test utilizes a single data loader, whereas the training process employs a separate data loader for each GPU. Consequently, the tight-loop read test establishes an upper bound on the data throughput that can be consumed by a single GPU.
To minimize AWS usage costs, each data loader was evaluated in a single test run for up to four epochs. The experimental results presented below compare the performance of the data loaders under varying latency conditions.
4.2.1 Tight-loop reading
As a baseline for comparing network data loaders, we measured the performance of NVIDIA DALI when reading images stored as TFRecords from the local filesystem, without performing image decoding. This configuration achieved a data throughput of 7.4 GB/s.
The p4d.24xlarge instance offers a total network bandwidth of 100 Gb/s; however, only half of this bandwidth is accessible through its public interface [^6] https://docs.aws.amazon.com/AWSEC2/latest/UserGuide/ec2-instance-network-bandwidth.html. Thus, the maximum raw bandwidth available for data transfers in our tests was limited to 50 Gb/s (6.25 GB/s). As shown in Fig. 3(a) and Tab. 3, our data loader nearly saturated the available bandwidth when reading from ScyllaDB in both local and medium-latency settings, achieving a throughput of approximately 6 GB/s in each case. In the high-latency setting, throughput decreased to around 4 GB/s.
In contrast, MosaicML SD demonstrated significantly lower performance, with throughput measured at 326 MB/s in the low-latency setting, 308 MB/s in the medium-latency setting, and 203 MB/s in the high-latency setting. tf.data service exhibited better performance than MosaicML SD in the low-latency configuration, achieving a throughput of 437 MB/s. However, its performance degraded substantially in higher-latency environments, with throughput dropping to 57 MB/s in the medium-latency setting and just 12 MB/s in the high-latency setting.
4.2.2 Training
For a data loader to be effective, its performance must integrate smoothly into the DL pipeline, ensuring that tensors are efficiently delivered to DL engines without introducing bottlenecks or delays. To evaluate this, we assessed the performance of data loaders within a standard image classification training workflow using the ResNet-50 architecture. Due to the differences in training performance between TensorFlow and PyTorch, we focused exclusively on one framework to ensure a fair comparison. Given that MosaicML SD demonstrated superior performance compared to tf.data service in medium- and high-latency settings in the previous evaluation, we chose to test it against our data loader in a training using PyTorch.
To establish a performance upper bound, we first performed training using a fixed input tensor, thereby eliminating the overhead associated with data loading and preprocessing. This setup enabled us to measure the maximum achievable data throughput during training. Specifically, we recorded the number of images processed per second on a single GPU and across all 8 GPUs during multi-GPU training. Our results indicate that a single NVIDIA A100 GPU consumes about 1450 images/s, while an 8-GPU configuration reaches 11200 images/s. Given the average ImageNet-1k training image size of 115 kB, the data loaders must sustain a steady throughput of approximately 1.3 GB/s to meet these requirements.
As demonstrated in Fig. 3(b) and Tab. 4, the MosaicML SD data loader is unable to sustain the throughput required to fully utilize all 8 GPUs, achieving 57%, 49%, and 33% of the target throughput under low, medium, and high-latency conditions, respectively. In contrast, our data loader achieves 94%, 95%, and 96% of the theoretical upper bound in these settings.
(a) Normalized tight-loop reading throughput.
(b) Normalized train reading throughput.
Figure 3: Normalized reading throughput at varying latencies. The red, dashed line in (a) shows the available network bandwidth.
Table 3: Tight-loop reading at varying latencies. The table shows average epoch data throughput and standard deviation (omitted when data is insufficient). Epoch throughput is calculated as the dataset size divided by epoch duration.
Data loader Throughput (MB/s)
Local Low-lat Med-lat Hi-lat
DALI TFRecord 7381 $\pm$ 108
MosaicML SD 326 $\pm$ 14 308 $\pm$ 15 203 $\pm$ 812
tf.data service 437 $\pm$ 512 57 $\pm$ 123 12 $\pm$ 123
Cassandra-DALI 6066 $\pm$ 147 5957 $\pm$ 115 4081 $\pm$ 337
Table 4: Pytorch ResNet-50 training at varying latencies. The table shows average epoch image throughput and standard deviation (omitted when data is insufficient). Epoch throughput is calculated as the dataset size divided by epoch duration.
Data loader Throughput (img/s)
Local Low-lat Med-lat Hi-lat
No I/O 11199 $\pm$ 312
MosaicML SD 6209 $\pm$ 89 5424 $\pm$ 735 3992 $\pm$ 512
Cassandra-DALI 10608 $\pm$ 113 10587 $\pm$ 300 10485 $\pm$ 98
4.3 Impact of prefetching and database choice on data loader performance
Finally, we conducted further tests to better investigate our data loader’s performance, focusing specifically on the impact of the proposed out-of-order prefetching optimization. Additionally, we analyzed how the choice of the underlying database – either Cassandra or ScyllaDB – affects data loading performance.
4.3.1 Impact of out-of-order, incremental prefetching
We evaluated the tight-loop reading performance under high-latency conditions, comparing results with and without our out-of-order, incremental prefetching optimization.
High-latency, high-bandwidth internet communications are prone to significant variability in TCP throughput. In fact these conditions exacerbate the effects of packet loss, as TCP congestion control mechanisms respond conservatively to retransmissions and recover slowly due to extended RTTs, resulting in reduced throughput [13, 8].
Figure 4 highlights the significant variance in batch loading times between in-order and out-of-order prefetching. In the in-order case (Figure 4(a)), when the prefetching queue is exhausted, the system experiences delays of up to several seconds while waiting for all transfers, including those over congested routes, to complete. This results in a cyclical pattern: once all transfers are completed, the queue is refilled, but it is quickly depleted again, triggering a new cycle. In contrast, the out-of-order approach (Figure 4(b)) maintains a highly consistent batch loading time, staying always below 30 ms after the initial transient period.
Figure 5 illustrates the throughput over time of each of the 32 TCP connections utilized by our data loader when employing the standard in-order prefetching mechanism. The throughput curves exhibit a strong correlation, highlighting that simultaneous transfers are constrained by the in-order batch assembly process, since the system must wait for the slowest transfer to finish before dispatching a batch to the DL pipeline and requesting a new batch from the database. As a result, the throughputs tend to converge and the aggregated throughput exhibits considerable fluctuations, ranging roughly from 300 MB/s to 1300 MB/s.
In contrast, relaxing this in-order constraint allows transfers to proceed independently, as shown in Figure 6. In this approach, batches are formed as soon as a sufficient number of images are available, irrespective of their originating connections. This optimization significantly enhances overall throughput, resulting in higher and more consistent performance, maintaining an average throughput of approximately 4 GB/s.
(a) In-order prefetching.
(b) Out-of-order prefetching.
Figure 4: Comparison of batch loading times in high-latency networks.
Figure 5: Transfer rate of 32 concurrent, high-latency TCP connections. In-order prefetching.
Figure 6: Transfer rate of 32 concurrent, high-latency TCP connections. Out-of-order prefetching.
4.3.2 Cassandra vs ScyllaDB
The tight-loop reading test under high-latency conditions was also performed using Cassandra as the storage backend for images, replacing ScyllaDB. As illustrated in Fig. 7, Cassandra achieved a throughput of 1.6 GB/s, significantly lower than the 4.0 GB/s observed with ScyllaDB, highlighting the superior performance of the latter. Notably, Cassandra exhibited a disk I/O rate considerably higher than its achieved data throughput (3.6 GB/s versus 1.6 GB/s), likely attributable to differences in its block-reading strategy compared to ScyllaDB.
Figure 7: Comparison of disk and network throughputs for Cassandra and ScyllaDB (tight-loop reading test).
5 Conclusion
The exponential growth in GPU computational power has enabled unprecedented advancements in deep learning, particularly for large-scale applications. However, the increasing discrepancy between processing throughput and data access latencies has introduced significant challenges in ensuring efficient data movement and storage management. Our proposed solution addresses these challenges by integrating scalable NoSQL databases with a high-performance data loader optimized for image-based DL tasks.
The key contributions of our work include a novel data loader designed to leverage state-of-the-art prefetching strategies, including out-of-order prefetching mechanisms that mitigate the impact of network variability and congestion, thereby maximizing data loading efficiency. By coupling data with metadata and employing a database-driven architecture, our implementation provides a scalable, flexible, and consistent solution for managing DL datasets. Through comprehensive evaluations under varying latency conditions, we demonstrated the effectiveness of our approach, achieving significant improvements in throughput and stability, particularly in high-latency environments. Comparative analyses further validate the robustness of our method against existing state-of-the-art data loading techniques.
The source code for our implementation is publicly available, providing a resource for further research and practical deployment in diverse DL scenarios.
Acknowledgment
This work was supported by the Italian Ministry of Health under the program grant H2ub (Hybrid Hub: Cellular and computation models, micro- and nano-technologies for personalized innovative therapies, project code T4-AN-10) and by the Regione Autonoma della Sardegna, Sardegna Ricerche, under the program grant XDATA.